# From tug-of-war to Brownian Boost: explicit ODE solutions for player-funded stochastic-differential games

### v3

@@ 5.2. Proving properties of the ODE pair
    $f_{\rho}(1,r)$ and $g_{\rho}(1,r)$ are bounded above by $e^{-2|r|+o(r)}$ as $r\to\infty$.
@@
    $\int_{\mathbb{R}}f_{\rho}(x,u)\,{\rm d}u=\int_{\mathbb{R}}g_{\rho}(x,u)\,{\rm d}u$.
Remark. The point $v=v(x)$ identified in the second part may be viewed as a battlefield location for $\big(f_{\rho}(x,\cdot),g_{\rho}(x,\cdot)\big)$, since the condition $S_{\rho}(x,v)=1$ is counterpart to $\phi_{k}$ being close to one for $k$ the battlefield index in the discrete case. This is not to say that gameplay at $v$ is uniquely influential. The battle occurs principally in a compact neighbourhood of $v$ (whose length depends on $\rho$).
Proof of Proposition 5.4(1,2). Write $J(r)=S_{\rho}(x,r)^{\rho}$ and integrate the differential equation in Lemma 5.1 on $[0,r]$. Since $J(0)=x^{\rho}$, we find that
$r=-\frac{1}{8\rho^{2}}\Big(H\big(J(r)\big)-H(x^{\rho})\Big)\,$,
@@
Rewriting the last display, $8\rho^{2}r=H\big(S_{\rho}(x,r)^{\rho}\big)-H(x^{\rho})$. Since $H(1)=0$, the unique solution $v=v(x)$ of $8\rho^{2}v=H(x^{\rho})$ (which is the value identified in the second part of the proposition) is that time for which $S_{\rho}(x,v)=1$ (as we seek to prove in that part). Now $r\to S_{\rho}(1,r)$ and $r\to S_{\rho}(x,v+r)$ solve the initial-value problem stated in Lemma 5.1. The uniqueness of the solution to this problem implied by this lemma shows that these two functions mapping $\mathbb{R}$ to $(0,\infty)$ are equal.
@@
$S_{\rho}(x,-r)=S_{\rho}(1,-r-v(x))=S_{\rho}(1,r+v(x))^{-1}=S_{\rho}(1,r-v(x^{-1}))^{-1}=S_{\rho}(x^{-1},r)\,$,
where the first and last equalities arise from the just obtained equality of functions applied for $x$ and $x^{-1}$. The second equality is an instance of $S_{\rho}(1,-r)=S_{\rho}(1,r)^{-1}$, while the third is due to $v(x^{-1})=-v(x)$, a fact seen from $8\rho^{2}v(x)=H(x^{\rho})=-H(x^{-\rho})=-8\rho^{2}v(x^{-1})$.
We use the representation (47) and (48) in the first and last equalities as we write, with $v=v(x)$,
$\frac{f_{\rho}(x,v+r)}{f_{\rho}(x,v)}$ $=$ $\exp\left\{2\int_{v}^{v+r}\Phi_{f}\big(S_{\rho}(x,u)\big)\,{\rm d}u\right\}\,=\,\exp\left\{2\int_{0}^{r}\Phi_{f}\big(S_{\rho}(x,v+s)\big)\,{\rm d}s\right\}$
$=$ $\exp\left\{2\int_{0}^{r}\Phi_{f}\big(S_{\rho}(1,s)\big)\,{\rm d}s\right\}=f_{\rho}(1,r)\,$,
the penultimate equality due to $S_{\rho}(1,r)=S_{\rho}(x,v+r)$. Thus we obtain the second part of the proposition in regard to $f$; the very similar argument for $g$ is omitted.
@@
$\Phi_{g}(1/s)=\Phi_{f}(s)\,$, (52)
since $\Phi_{g}(1/s)=1-\frac{2\big((1-\rho)s^{-\rho}+s^{-2\rho}\big)}{(1+s^{-\rho})^{2}}=1-\frac{2\big((1-\rho)s^{\rho}+1\big)}{(1+s^{\rho})^{2}}=\Phi_{f}(s)$. Using again the expressions (47) and (48),
$\log g_{\rho}(1,-r)$ $=$ $-2\int_{0}^{-r}\Phi_{g}\big(S_{\rho}(1,u)\big)\,{\rm d}u=2\int_{0}^{r}\Phi_{g}\big(S_{\rho}(1,-w)\big)\,{\rm d}w$
$=$ $2\int_{0}^{r}\Phi_{f}\big(S_{\rho}(1,w)\big)\,dw=\log f_{\rho}(1,r)\,$,
where $\Phi_{g}\big(S_{\rho}(1,-w)\big)=\Phi_{f}\big(S_{\rho}(1,w)\big)$ is due to Proposition 5.4(1) and (52). Exponentiating, we obtain the sought statement.
(4). As the solution to the differential equation in Definition 1.2, $S_{\rho}(x,u)>0$ is readily seen to converge to zero and infinity in the respective limits of large positive and negative $u$. So
$\Phi_{f}(S_{\rho}(x,u))\to\begin{cases}1&\textrm{as}\,\,\,u\to-\infty\\ -1&\textrm{as}\,\,\,u\to\infty\end{cases}\,\,\,\,\textrm{and}\,\,\,\,\Phi_{g}(S_{\rho}(x,u))\to\begin{cases}-1&\textrm{as}\,\,\,u\to-\infty\\ 1&\textrm{as}\,\,\,u\to\infty\end{cases}\,.$
Note that the convention $\int_{a}^{b}f=-\int_{b}^{a}f$ is in force as we interpret (47) and (48). We see that $|r|^{-1}\log f_{\rho}(1,r)$ and $|r|^{-1}\log g_{\rho}(1,r)$ converge to $2$, as $r$ tends to both minus and plus infinity. This yields the sought statement.
(5). First note the special case when $x=1$: $\int_{\mathbb{R}}f_{\rho}(1,u)\,{\rm d}u=\int_{\mathbb{R}}g_{\rho}(1,u)\,{\rm d}u$. This is due to the symmetry and integrability offered by the preceding two parts.
@@
$\frac{1}{f_{\rho}(x,v)}\int_{\mathbb{R}}f_{\rho}(x,u)\,{\rm d}u=\int_{\mathbb{R}}f_{\rho}(1,u)\,{\rm d}u=\int_{\mathbb{R}}g_{\rho}(1,u)\,{\rm d}u=\frac{1}{g_{\rho}(x,v)}\int_{\mathbb{R}}g_{\rho}(x,u)\,{\rm d}u\,.$
Here, Proposition 5.4(2) furnishes $v=v(x)$, and the other displayed equalities are obtained by integrating the identities in this result over $\mathbb{R}$.
As noted in Theorem 1.3, $g_{\rho}(x,\cdot)=f_{\rho}(x,\cdot)S_{\rho}(x,\cdot)$. Hence,
$\frac{\int_{\mathbb{R}}g_{\rho}(x,u)\,{\rm d}u}{\int_{\mathbb{R}}f_{\rho}(x,u)\,{\rm d}u}\;=\;\frac{g_{\rho}(x,v)}{f_{\rho}(x,v)}=S_{\rho}(x,v)\,.$
But by Proposition 5.4(2), $S_{\rho}(x,v)=1$: so the integrals are equal. ∎
We now prove the high $|u|$ asymptotics of $f_{\rho}(1,u)$ and $g_{\rho}(1,u)$, thereby refining Proposition 5.4(4), and concomitant results for the stake functions $a_{\rho}(1,u)$ and $b_{\rho}(1,u)$.
Proof of Proposition 1.5. As remarked after Lemma 5.1, $J(u)=S_{\rho}(1,u)^{\rho}$ satisfies $J(u)\sim\frac{1}{8\rho^{2}u}$ for $u\gg 0$. In the representation (47) and (48), $J$ enters in the role of $S$ in the functions $\Phi_{f}$ and $\Phi_{g}$; for small $J$, $\Phi_{f}=-1+2(1+\rho)J+O(J^{2})$ and $\Phi_{g}=1-2(1-\rho)J+O(J^{2})$. Using $\int_{0}^{u}J(w)\,{\rm d}w\sim\frac{1}{8\rho^{2}}\log u$ and $\int_{0}^{u}J(w)^{2}\,{\rm d}w=O(1)$, we take $r=u$ in these representations to obtain
@@
By Definition 1.4, $S_{\rho}(x,u)=g_{\rho}(x,u)/f_{\rho}(x,u)$ (from Theorem 1.3), and $J(u)=S_{\rho}(1,u)^{\rho}$, we see that
@@
6. The high-noise limit
Extending a specification (with $r=\pm\infty$) made after Definition 1.4 in regard to $\mathrm{BB}(\rho)$ ODE pair solutions $\big(f_{\rho}(x,\cdot),g_{\rho}(x,\cdot)\big)$, we set $m_{\rho}(x,r)=\int_{-\infty}^{r}f_{\rho}(x,u)\,{\rm d}u$ and $n_{\rho}(x,r)=\int_{r}^{\infty}g_{\rho}(x,u)\,{\rm d}u$ for $r\in\mathbb{R}$. In this way, ABMN($\kappa,\rho$) elements have Brownian Boost counterparts $(a_{\rho},b_{\rho},m_{\rho},n_{\rho})$. Here, we study ABMN($\kappa,\rho$) elements in the limit of low $\kappa$, showing how they converge to their $\mathrm{BB}(\rho)$ counterpart, and reaching such conclusions as Theorem 1.18.
@@ 6.1. Two routes to Brownian Boost
In this subsection, we present a four-part proposition concerning ABMN elements whose first two parts offer simple and useful stake formulas and whose latter parts permit us to discuss competing routes to our analysis of $\rho$-Brownian Boost. After the discussion and proof, we will signpost the structure of Section 6.
@@
The equations in Proposition 6.1(3,4) are discrete counterparts to the $\rho$-Brownian Boost ODE pair (2), the pairs’ respective elements identified under the correspondence of $m_{i,i+1}$ with $m^{\prime}=f$ and $n_{i,i-1}$ with $n^{\prime}=-g^{\prime}$. Indeed, suppose that we permit the comparisons $\kappa^{-1}m_{\kappa^{-1}u-1,\kappa^{-1}u}=f(u)+O(\kappa)$ and $\kappa^{-1}n_{\kappa^{-1}u,\kappa^{-1}u-1}=-g(u)+O(\kappa)$ and their corollaries $\Delta m_{\kappa^{-1}u}=\kappa^{2}f^{\prime}(u)+O(\kappa^{3})$, $\Delta n_{\kappa^{-1}u}=-\kappa^{2}g^{\prime}(x)+O(\kappa^{3})$, $M_{\kappa^{-1}u}=2\,m_{\kappa^{-1}u-1,\kappa^{-1}u}+O(\kappa^{2})$ and $N_{\kappa^{-1}u}=2\,n_{\kappa^{-1}u,\kappa^{-1}u-1}+O(\kappa^{2})$. Then on dividing the Proposition 6.1(3,4) equations by $2^{2\rho}\kappa^{2(1+\rho)}$, we would learn that $f$ and $g$ satisfy the ODE pair (2) up to an $O(\kappa)$ error that must vanish since $\kappa>0$ may tend to zero. Suitably elaborated, such an approach lead to a rigorous discrete counterpart to the analysis of Brownian Boost offered in Section 5.1 wherein (2) was heuristically derived.
@@
Proof of Proposition 6.1(1,2). Use of the shorthand $_{i,j}=*_{j}-*_{i}$ for $\in\{m,n\}$ continues. Analogous to $a=\frac{a^{\rho}-b^{\rho}}{a^{\rho}+b^{\rho}}m^{\prime}+m^{\prime\prime}/2$ and to $b=\frac{a^{\rho}-b^{\rho}}{a^{\rho}+b^{\rho}}n^{\prime}+n^{\prime\prime}/2$ in (41) and (42) are the equations
@@
We seek a counterpart to (41). Rearranging the above gives
$a_{i}\,=\,-\frac{b_{i}^{\rho}}{a_{i}^{\rho}+b_{i}^{\rho}}\kappa\cdot m_{i-1,i+1}\,+\,\kappa\cdot m_{i,i+1}\,+\,\frac{1-\kappa}{2}\Delta_{i}m$ (53)
@@
$b_{i}\,=\,-\frac{a_{i}^{\rho}}{a_{i}^{\rho}+b_{i}^{\rho}}\kappa\cdot n_{i+1,i-1}\,-\,\kappa\cdot n_{i,i-1}\,+\,\frac{1-\kappa}{2}\Delta_{i}n\,.$ (54)
@@
$a_{i}=\frac{\kappa\rho M_{i}^{1+\rho}N_{i}^{\rho}}{(M_{i}^{\rho}+N_{i}^{\rho})^{2}}\qquad\textrm{and}\qquad b_{i}=\frac{\kappa\rho M_{i}^{\rho}N_{i}^{1+\rho}}{(M_{i}^{\rho}+N_{i}^{\rho})^{2}}\,$, (55)
@@
(3,4). Returning to (53) and (54) with the expressions (55), and multiplying both of the resulting equations by $\big(M_{i}^{\rho}+N_{i}^{\rho}\big)^{2}$, we obtain
$\kappa\rho M_{i}^{1+\rho}N_{i}^{\rho}\,=\,\kappa M_{i}^{\rho}\big(M_{i}^{\rho}+N_{i}^{\rho}\big)m_{i,i+1}\,-\,\kappa N_{i}^{\rho}\big(M_{i}^{\rho}+N_{i}^{\rho}\big)m_{i-1,i}\,+\,\tfrac{1-\kappa}{2}\cdot\big(M_{i}^{\rho}+N_{i}^{\rho}\big)^{2}\Delta_{i}m$ (56)
@@
$\kappa\rho M_{i}^{\rho}N_{i}^{1+\rho}\,=\,\kappa N_{i}^{\rho}\big(M_{i}^{\rho}+N_{i}^{\rho}\big)n_{i,i-1}\,-\,\kappa M_{i}^{\rho}\big(M_{i}^{\rho}+N_{i}^{\rho}\big)n_{i+1,i}\,+\,\tfrac{1-\kappa}{2}\cdot\big(M_{i}^{\rho}+N_{i}^{\rho}\big)^{2}\Delta_{i}n\,.$ (57)
@@
Taking the difference of these equations, we may substitute the outcome into (56), thereby finding that
$\kappa\rho M_{i}^{1+\rho}N_{i}^{\rho}\,=\,\tfrac{\kappa}{2}\cdot M_{i}\big(M_{i}^{2\rho}-N_{i}^{2\rho}\big)\,+\,\tfrac{1}{2}\cdot\big(M_{i}^{\rho}+N_{i}^{\rho}\big)^{2}\Delta_{i}m\,.$ (58)
@@
which substituted into (57) yield
$\kappa\rho M_{i}^{\rho}N_{i}^{1+\rho}\,=\,\tfrac{\kappa}{2}\cdot N_{i}\big(N_{i}^{2\rho}-M_{i}^{2\rho}\big)\,+\,\tfrac{1}{2}\cdot\big(M_{i}^{\rho}+N_{i}^{\rho}\big)^{2}\Delta_{i}n\,$, (59)
@@
There are five further subsections. In the next three, we compare one-step of the application $s:\phi_{0}\mapsto\phi_{1}$ to a suitably short passage along the $S_{\rho}$-flow; infer how a $\kappa^{-1}$-speeding of time leads to a description via this flow; and prove as a consequence the stake-asymptotic Theorem 1.18(1). In the two further subsections, we prove Theorem 1.22 on the approach of $\lambda_{\rm max}$ to one; and derive the scaled gameplay Theorem 1.18(2).

### v4

@@ 5.2. Proving properties of the ODE pair
    $f_{\rho}(1,u)$ and $g_{\rho}(1,u)$ are bounded above by $e^{-2|u|+o(|u|)}$ as $|u|\to\infty$.
@@
    $\int_{\mathbb{R}}f_{\rho}(\xi,r)\,{\rm d}r=\int_{\mathbb{R}}g_{\rho}(\xi,r)\,{\rm d}r$.
Remark. In signposting Brownian Boost in Subsection 1.1.2, we indicated that TINE are indexed by a battlefield value $v\in\mathbb{R}$, the equilibrium profile for this general value being obtained from the zero-indexed profile by a shift of $v$. It is Proposition 5.4(2) that identifies the value $v=v(\xi)$. Since $\xi\mapsto v$ is a bijection $(0,\infty)\to\mathbb{R}$, the flow index and battlefield parameterisations are equivalent. However, translation invariance of equilibrium profiles is more naturally expressed via the game theoretic interpretation, using the standard normalization for solutions specified in Definition 1.17, rather than by means of the default normalization we are presently using in this analytic discussion. We will elaborate at the end of the chapter.
Proof of Proposition 5.4(1,2). Write $J(u)=S_{\rho}(\xi,u)^{\rho}$ and integrate the differential equation in Lemma 5.1 on $[0,u]$. Since $J(0)=\xi^{\rho}$, we find that
$u=-\frac{1}{8\rho^{2}}\Big(H\big(J(u)\big)-H(\xi^{\rho})\Big)\,$,
@@
The function $H$ is an increasing bijection from $(0,\infty)$ to $\mathbb{R}$; thus, so is $\xi\mapsto H(\xi^{\rho})$, and also $\xi\mapsto v(\xi)$, as we claim in the second part.
@@
Rewriting the last display, $8\rho^{2}u=H\big(S_{\rho}(\xi,u)^{\rho}\big)-H(\xi^{\rho})$. Since $H(1)=0$, the unique solution $v=v(\xi)$ of $8\rho^{2}v=H(\xi^{\rho})$ (which is the value identified in the second part of the proposition) is that time for which $S_{\rho}(\xi,v)=1$ (as we seek to prove in that part). Now $u\to S_{\rho}(1,u)$ and $u\to S_{\rho}(\xi,v+u)$ solve the initial-value problem stated in Lemma 5.1. The uniqueness of the solution to this problem implied by this lemma shows that these two functions mapping $\mathbb{R}$ to $(0,\infty)$ are equal.
@@
$S_{\rho}(\xi,-u)=S_{\rho}(1,-u-v(\xi))=S_{\rho}(1,u+v(\xi))^{-1}=S_{\rho}(1,u-v(\xi^{-1}))^{-1}=S_{\rho}(\xi^{-1},u)^{-1}\,$,
where the first and last equalities arise from the just obtained equality of functions applied for $\xi$ and $\xi^{-1}$. The second equality is an instance of $S_{\rho}(1,-u)=S_{\rho}(1,u)^{-1}$, while the third is due to $v(\xi^{-1})=-v(\xi)$, a fact seen from $8\rho^{2}v(\xi)=H(\xi^{\rho})=-H(\xi^{-\rho})=-8\rho^{2}v(\xi^{-1})$.
We use the representation (5.7) and (5.8) in the first and last equalities as we write, with $v=v(\xi)$,
$\frac{f_{\rho}(\xi,v+u)}{f_{\rho}(\xi,v)}$ $=$ $\exp\left\{2\int_{v}^{v+u}\Phi_{f}\big(S_{\rho}(\xi,r)\big)\,{\rm d}r\right\}\,=\,\exp\left\{2\int_{0}^{u}\Phi_{f}\big(S_{\rho}(\xi,v+r)\big)\,{\rm d}r\right\}$
$=$ $\exp\left\{2\int_{0}^{u}\Phi_{f}\big(S_{\rho}(1,r)\big)\,{\rm d}r\right\}=f_{\rho}(1,u)\,$,
the penultimate equality due to $S_{\rho}(1,r)=S_{\rho}(\xi,v+r)$. Thus we obtain the second part of the proposition in regard to $f$; the very similar argument for $g$ is omitted.
@@
$\Phi_{g}(1/s)=\Phi_{f}(s)\,$, (5.12)
since $\Phi_{g}(1/s)=1-\frac{2\big((1-\rho)s^{-\rho}+s^{-2\rho}\big)}{(1+s^{-\rho})^{2}}=1-\frac{2\big((1-\rho)s^{\rho}+1\big)}{(1+s^{\rho})^{2}}=\Phi_{f}(s)$. Using again the expressions (5.7) and (5.8),
$\log g_{\rho}(1,-u)$ $=$ $-2\int_{0}^{-u}\Phi_{g}\big(S_{\rho}(1,r)\big)\,{\rm d}r=2\int_{0}^{u}\Phi_{g}\big(S_{\rho}(1,-r)\big)\,{\rm d}r$
$=$ $2\int_{0}^{u}\Phi_{f}\big(S_{\rho}(1,r)\big)\,{\rm d}r=\log f_{\rho}(1,u)\,$,
where $\Phi_{g}\big(S_{\rho}(1,-r)\big)=\Phi_{f}\big(S_{\rho}(1,r)\big)$ is due to Proposition 5.4(1) and (5.12). Exponentiating, we obtain the sought statement.
(4). As the solution to the differential equation in Definition 1.14, $S_{\rho}(\xi,u)>0$ is readily seen to converge to zero and infinity in the respective limits of large positive and negative $u$. So
$\Phi_{f}(S_{\rho}(\xi,u))\to\begin{cases}1&\textrm{as}\,\,\,u\to-\infty\\ -1&\textrm{as}\,\,\,u\to\infty\end{cases}\,\,\,\,\textrm{and}\,\,\,\,\Phi_{g}(S_{\rho}(\xi,u))\to\begin{cases}-1&\textrm{as}\,\,\,u\to-\infty\\ 1&\textrm{as}\,\,\,u\to\infty\end{cases}\,.$
Note that the convention $\int_{a}^{b}f=-\int_{b}^{a}f$ is in force as we interpret (5.7) and (5.8). We see that $|u|^{-1}\log f_{\rho}(1,u)$ and $|u|^{-1}\log g_{\rho}(1,u)$ converge to $2$, as $u$ tends to both minus and plus infinity. This yields the sought statement.
(5). First note the special case when $\xi=1$: $\int_{\mathbb{R}}f_{\rho}(1,r)\,{\rm d}r=\int_{\mathbb{R}}g_{\rho}(1,r)\,{\rm d}r$. This is due to the symmetry and integrability offered by the preceding two parts.
@@
$\frac{1}{f_{\rho}(\xi,v)}\int_{\mathbb{R}}f_{\rho}(\xi,r)\,{\rm d}r=\int_{\mathbb{R}}f_{\rho}(1,r)\,{\rm d}r=\int_{\mathbb{R}}g_{\rho}(1,r)\,{\rm d}r=\frac{1}{g_{\rho}(\xi,v)}\int_{\mathbb{R}}g_{\rho}(\xi,r)\,{\rm d}r\,.$
Here, Proposition 5.4(2) furnishes $v=v(\xi)$, and the other displayed equalities are obtained by integrating the identities in this result over $\mathbb{R}$.
As noted in Theorem 1.15, $g_{\rho}(\xi,\bullet)=f_{\rho}(\xi,\bullet)S_{\rho}(\xi,\bullet)$. Hence,
$\frac{\int_{\mathbb{R}}g_{\rho}(\xi,r)\,{\rm d}r}{\int_{\mathbb{R}}f_{\rho}(\xi,r)\,{\rm d}r}\;=\;\frac{g_{\rho}(\xi,v)}{f_{\rho}(\xi,v)}=S_{\rho}(\xi,v)\,.$
But by Proposition 5.4(2), $S_{\rho}(\xi,v)=1$: so the integrals are equal. ∎
We turn to proving the high $|u|$ asymptotics of $f_{\rho}(1,u)$ and $g_{\rho}(1,u)$, thereby refining Proposition 5.4(4), and concomitant results for the stake functions $a_{\rho}(1,u)$ and $b_{\rho}(1,u)$.
Proof of Proposition 1.19. As remarked after Lemma 5.1, $J(u)=S_{\rho}(1,u)^{\rho}$ satisfies $J(u)\sim\frac{1}{8\rho^{2}u}$ for $u\gg 0$. In the representations (5.7) and (5.8), $J$ enters in the role of $S$ in the functions $\Phi_{f}$ and $\Phi_{g}$; for small $J$, $\Phi_{f}=-1+2(1+\rho)J+O(J^{2})$ and $\Phi_{g}=1-2(1-\rho)J+O(J^{2})$. Using the high-$u$ asymptotics $\int_{0}^{u}J(r)\,{\rm d}r\sim\frac{1}{8\rho^{2}}\log u$ and $\int_{0}^{u}J(r)^{2}\,{\rm d}r=O(1)$ in these representations, we obtain
@@
By Definition 1.16, $S_{\rho}(\xi,u)=g_{\rho}(\xi,u)/f_{\rho}(\xi,u)$ (from Theorem 1.15), and $J(u)=S_{\rho}(1,u)^{\rho}$, we see that
@@
We end the chapter by offering the explanation, promised after Proposition 5.4, that concerns the translation invariance of $\mathrm{BB}(\rho)$-equilibrium profiles and the indexing role for the battlefield value $v=v(\xi)$.
Noting that $v(1)=0$, we see from the the relation $S_{\rho}(\xi,v(\xi))=1$ that default solutions satisfy
$\big(f_{\rho}(\xi,\bullet),g_{\rho}(\xi,\bullet)\big)=C(\xi)\,\big(f_{\rho}(1,\bullet-v(\xi)),\,g_{\rho}(1,\bullet-v(\xi))\big)$
for some positive factor $C(\xi)$. In the present context, however, the suitable normalization is the standard one, from Definition 1.17, in accordance with our convention that Maxine’s winning receipt in $\mathrm{BB}(\rho)$ equals one. The counterpart to the preceding display for standard solutions has $C(\xi)$ identically equal to one, for it is this choice that ensures the property that $\int_{\mathbb{R}}f_{\rho}(\xi,r){\rm d}r$ is independent of $\xi\in(0,\infty)$ which is required in the standard normalization. Note that the formulas in Definition 1.16 for equilibrium profiles are homogeneous of degree one in the inputs $(f,g)$. Writing $(a,b)$ for the profile thus associated to the value $\xi=1$, we thus confirm that the profile indexed by $\xi\in(0,\infty)$ takes the form $\big(a(\bullet-v),\,b(\bullet-v)\big)$, with $v=v(\xi)$.
Chapter 6 The fine-mesh high-noise limit
When $\mathrm{TLP}(\kappa,\rho)$ is viewed in the fine-mesh, high-noise limit $\kappa\searrow 0$, space is squeezed (by a factor of $\kappa$), time is sped up (by a factor of $\kappa^{-2}$), and stakes are revalued (again by a factor of $\kappa^{-2}$). In our treatment, Brownian Boost is given a rigorous interpretation as the scaling limit of the Trail of Lost Pennies, with its time-invariant equilibria governed by the $\rho$-Brownian Boost ODE pair. This chapter is devoted to obtaining the scaled description and to drawing out some of its consequences.
@@ 6.1. Two routes to Brownian Boost
In this section, we present a four-part proposition concerning ABMN elements whose first two parts offer simple and useful stake formulas and whose latter parts permit us to discuss competing routes to our analysis of $\rho$-Brownian Boost. After the discussion and proof, we will signpost the structure of Chapter 6.
@@
The equations in Proposition 6.1(3,4) are discrete counterparts to the $\rho$-Brownian Boost ODE pair (1.13), the pairs’ respective elements identified under the correspondence of $m_{i,i+1}$ with $m^{\prime}=f$ and $n_{i,i-1}$ with $n^{\prime}=-g^{\prime}$. Indeed, suppose that we permit the comparisons $\kappa^{-1}m_{\kappa^{-1}u-1,\kappa^{-1}u}=f(u)+O(\kappa)$ and $\kappa^{-1}n_{\kappa^{-1}u,\kappa^{-1}u-1}=-g(u)+O(\kappa)$ and their corollaries $\Delta m_{\kappa^{-1}u}=\kappa^{2}f^{\prime}(u)+O(\kappa^{3})$, $\Delta n_{\kappa^{-1}u}=-\kappa^{2}g^{\prime}(x)+O(\kappa^{3})$, $M_{\kappa^{-1}u}=2\,m_{\kappa^{-1}u-1,\kappa^{-1}u}+O(\kappa^{2})$ and $N_{\kappa^{-1}u}=2\,n_{\kappa^{-1}u,\kappa^{-1}u-1}+O(\kappa^{2})$. Then on dividing the Proposition 6.1(3,4) equations by $2^{2\rho}\kappa^{2(1+\rho)}$, we would learn that $f$ and $g$ satisfy the ODE pair (1.13) up to an $O(\kappa)$ error that must vanish since $\kappa>0$ may tend to zero. Suitably elaborated, such an approach would lead to a rigorous discrete counterpart to the analysis of Brownian Boost offered in Section 5.1 wherein (1.13) was heuristically derived.
@@
Proof of Proposition 6.1(1,2). Use of the shorthand $_{i,j}=*_{j}-*_{i}$ for $\in\{m,n\}$ continues. Analogous to $a=\frac{a^{\rho}-b^{\rho}}{a^{\rho}+b^{\rho}}m^{\prime}+m^{\prime\prime}/2$ and to $b=\frac{a^{\rho}-b^{\rho}}{a^{\rho}+b^{\rho}}n^{\prime}+n^{\prime\prime}/2$ in (5.1) and (5.2) are the equations
@@
We seek a counterpart to (5.1). Rearranging the above gives
$a_{i}\,=\,-\frac{b_{i}^{\rho}}{a_{i}^{\rho}+b_{i}^{\rho}}\kappa\cdot m_{i-1,i+1}\,+\,\kappa\cdot m_{i,i+1}\,+\,\frac{1-\kappa}{2}\Delta_{i}m$ (6.1)
@@
$b_{i}\,=\,-\frac{a_{i}^{\rho}}{a_{i}^{\rho}+b_{i}^{\rho}}\kappa\cdot n_{i+1,i-1}\,-\,\kappa\cdot n_{i,i-1}\,+\,\frac{1-\kappa}{2}\Delta_{i}n\,.$ (6.2)
@@
$a_{i}=\frac{\kappa\rho M_{i}^{1+\rho}N_{i}^{\rho}}{(M_{i}^{\rho}+N_{i}^{\rho})^{2}}\qquad\textrm{and}\qquad b_{i}=\frac{\kappa\rho M_{i}^{\rho}N_{i}^{1+\rho}}{(M_{i}^{\rho}+N_{i}^{\rho})^{2}}\,$, (6.3)
@@
(3,4). Returning to (6.1) and (6.2) with the expressions (6.3), and multiplying both of the resulting equations by $\big(M_{i}^{\rho}+N_{i}^{\rho}\big)^{2}$, we obtain
$\kappa\rho M_{i}^{1+\rho}N_{i}^{\rho}\,=\,\kappa M_{i}^{\rho}\big(M_{i}^{\rho}+N_{i}^{\rho}\big)m_{i,i+1}\,-\,\kappa N_{i}^{\rho}\big(M_{i}^{\rho}+N_{i}^{\rho}\big)m_{i-1,i}\,+\,\tfrac{1-\kappa}{2}\cdot\big(M_{i}^{\rho}+N_{i}^{\rho}\big)^{2}\Delta_{i}m$ (6.4)
@@
$\kappa\rho M_{i}^{\rho}N_{i}^{1+\rho}\,=\,\kappa N_{i}^{\rho}\big(M_{i}^{\rho}+N_{i}^{\rho}\big)n_{i,i-1}\,-\,\kappa M_{i}^{\rho}\big(M_{i}^{\rho}+N_{i}^{\rho}\big)n_{i+1,i}\,+\,\tfrac{1-\kappa}{2}\cdot\big(M_{i}^{\rho}+N_{i}^{\rho}\big)^{2}\Delta_{i}n\,.$ (6.5)
@@
Taking the difference of these equations, we may substitute the outcome into (6.4), thereby finding that
$\kappa\rho M_{i}^{1+\rho}N_{i}^{\rho}\,=\,\tfrac{\kappa}{2}\cdot M_{i}\big(M_{i}^{2\rho}-N_{i}^{2\rho}\big)\,+\,\tfrac{1}{2}\cdot\big(M_{i}^{\rho}+N_{i}^{\rho}\big)^{2}\Delta_{i}m\,.$ (6.6)
@@
which substituted into (6.5) yield
$\kappa\rho M_{i}^{\rho}N_{i}^{1+\rho}\,=\,\tfrac{\kappa}{2}\cdot N_{i}\big(N_{i}^{2\rho}-M_{i}^{2\rho}\big)\,+\,\tfrac{1}{2}\cdot\big(M_{i}^{\rho}+N_{i}^{\rho}\big)^{2}\Delta_{i}n\,$, (6.7)
@@
There are five further sections in this chapter. In the next three, we compare one-step of the application $s:\phi_{0}\mapsto\phi_{1}$ to a suitably short passage along the $S_{\rho}$-flow; infer how a $\kappa^{-1}$-speeding of time leads to a description via this flow; and prove as a consequence the stake-asymptotic Theorem 1.18(1). In the two further sections, we prove Theorem 1.23 on the approach of $\lambda_{\rm max}$ to one; and derive the scaled gameplay Theorem 1.18(2).